OpenStudy (anonymous):

find the value of ƒ ^-1 (4)

Vocaloid (vocaloid):

you would need to know the original function

Vocaloid (vocaloid):

is there a graph? a function?

OpenStudy (spring98):

\[(4) 1/f \] is the answer because you have to use the rule \[\:a^{-1}=\frac{1}{a}\] \[f^{-1}=\frac{1}{f}\] = \[=\left(4\right)\frac{1}{f}\]

OpenStudy (spring98):

I hope this helped!! :}

OpenStudy (solomonzelman):

I think that \(\large\color{#996633}{\displaystyle f^{-1}(4)}\) is a notation for an inverse-function of some function \(\large\color{#996633}{\displaystyle f(x)}\), evaluated at \(\large\color{#996633}{\displaystyle x=4}\).